• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name jpchen

    Log entry time 15:51:28 on April 1,2005

    Entry number 142980

    This entry is a followup to: 142963

    keyword=day shift summary continue

    4/1/05 dau shift summary continue:

    10:30 still smooth running.
    11:50 Ran raised HV for N3-19R by 40 V, whihc is during run 3128
    13:15 MCC took beam away to tune, try to improve beam quality.
    Compton photon detector is turned off. Elelctron detector is on and used to check background/beam quality (~300 is good, ~3000 is high).
    13:50 beam back. Compton electron rate is still high ~3000. MCC is
    still looking into it.
    14:20 MCC tuned beam and made the Compton electron rate down to
    a few hundred Hz.

    run 3127, 10:20-11:26, 300K
    run 3128, 11:28-12:38, 300K, HV N3-19R raised by 40V during run.
    run 3129, 12:39-13:16, 171K. MCC took beam away to tune beam quality.
    run 3130, 13:55-15:05, 300K.
    run 3131, 15:05-, still ongoing at the end of shift.